# references.bib

### v2

@@ 2 Modeling with ODEs
Our point estimate of $\hat{\theta}$ gives us the value of $\theta$ that means that $T(t)$ is the best match to the data, but this point estimate does not provide any indication of uncertainty in our estimate of $\theta$. To address our second question (above) we will work with the normalised log-likelihood function
@@
This process of holding one interest parameter constant and optimising out the remaining nuisance parameters by maximising the log-likelihood function can be repeated for all components of $\theta$. In Figure 1 we only have two parameters so we have two univariate profile likelihood functions. We can obtain the second profile likelihood by treating $k$ as the interest parameter across a uniform disretisation $k=0.02,0.025,0.03,\ldots,0.10$ /minute, and for each value of $k$ along the uniform discretisation, $k^{\dagger}$, we identify the value of $T_{a}$ that maximises $\bar{\ell}$ along the horizontal line where $k=k^{\dagger}$. Repeating this process for all values of $k$ along the uniform discretisation of $k$ generates a univariate profile likelihood function for the interest parameter $k$.
For our simple problem with two unknown parameters the process of fixing one interest parameter and optimising out the other nuisance parameter has a simple graphical interpretation that is unclear for higher dimensional problems with three or more parameters. Therefore, in general we use numerical optimization to construct univariate profile likelihood functions since this approach can be used regardless of the number of unknown parameters. Results in Figure 2 show the univariate profile likelihood functions for $T_{a}$ and $k$ obtained using the same numerical optimisation algorithm that we used to estimate the MLE. Both univariate profile likelihood functions involve a single, well-defined peak at the MLE indicating that these parameters are identifiable.
@@ 3 Modeling with PDEs
where $D>0$ [L2/T] is the diffusion coefficient, and $v$ [L/T] is the advective velocity. This mathematical model is widely used in various physical, chemical and biological applications, such as the study of the dispersion of dissolved solutes (e.g. nutrients, pollutants) in porous media [Zheng2002]. We will consider the solution of (4) for the initial condition $u(x,0)=u_{b}+u_{0}$ for $x<|h|$, and $u(x,0)=u_{b}$ for $x>|h|$. We interpret $u_{b}\geq 0$ as a uniform background concentration of $u$, and we are particularly interested in the spatial spreading of $u$ that arises when an additional amount of $u$ is placed uniformly within distance $h>0$ from the origin, giving $u(x,0)=u_{b}+u_{0}$ for $x<|h|$, where $u_{0}>0$. This initial condition is plotted in Figure 3 where we have a background concentration $u_{b}=1$, and locally within the interval $-50<x<50$ we have $u(x,0)=2$. The solution of the mathematical model describes how this additional solute within $-50<x<50$ undergoes combined advection and diffusive transport as a function of position $x$ and time $t$.
@@
Given our log-likelihood function and our estimate of $\hat{\theta}$, we can work with a normalised log-likelihood function and repeat the construction of the four univariate profile likelihood functions in exactly the same way as we did for the additive Gaussian noise model. The univariate profile likelihood functions in Figure 4 indicate that the four parameters are practically identifiable. As before, we can sample the log-likelihood function to obtain $M$ parameter samples $\theta$ with $\bar{\ell}\geq\bar{\ell}^{*}$ and use these $M$ solutions of (4) to construct the 95% prediction interval given in Figure 4. For the multiplicative noise model we see that the prediction intervals have the attractive property that the upper and lower bounds are always non-negative. This is very different to working with an additive Gaussian noise model since this procedure can lead to negative prediction intervals which is not physically realistic.
@@ 5 Extensions and general remarks
In terms of making model predictions, we always used a very simple rejection sampling method to find $M$ samples of $\theta$ where the log-likelihood function is great than some asymptotic threshold. We chose to use rejection sampling because it is both simple to implement and interpret, but other approaches are possible. For example, we could have simply used a uniform discretisation of the log-likelihood function and evaluated $\bar{\ell}$ across a uniform mesh to propagate the parameter confidence set through to examine model predictions. Both approaches carry advantages and disadvantages and give very similar results provided that $M$ is chosen to be sufficiently large when using rejection sampling, and that the uniform mesh is taken to be sufficiently dense when working with a gridded log-likelihood function.

### v3

@@ 2 Modeling with ODEs
Our point estimate of $\hat{\theta}$ gives us the value of $\theta$ that means that $T(t)$ is the best match to the data, but this point estimate does not provide any indication of uncertainty in our estimate of $\theta$. To address our second question (above) we will work with the normalized log-likelihood function
@@
This process of holding one interest parameter constant and optimizing out the remaining nuisance parameters by maximising the log-likelihood function can be repeated for all components of $\theta$. In Figure 1 we only have two parameters so we have two univariate profile likelihood functions. We can obtain the second profile likelihood by treating $k$ as the interest parameter across a uniform disretisation $k=0.02,0.025,0.03,\ldots,0.10$ /minute, and for each value of $k$ along the uniform discretisation, $k^{\dagger}$, we identify the value of $T_{a}$ that maximises $\bar{\ell}$ along the horizontal line where $k=k^{\dagger}$. Repeating this process for all values of $k$ along the uniform discretisation of $k$ generates a univariate profile likelihood function for the interest parameter $k$.
For our simple problem with two unknown parameters the process of fixing one interest parameter and optimizing out the other nuisance parameter has a simple graphical interpretation that is unclear for higher dimensional problems with three or more parameters. Therefore, in general we use numerical optimization to construct univariate profile likelihood functions since this approach can be used regardless of the number of unknown parameters. Results in Figure 2 show the univariate profile likelihood functions for $T_{a}$ and $k$ obtained using the same numerical optimisation algorithm that we used to estimate the MLE. Both univariate profile likelihood functions involve a single, well-defined peak at the MLE indicating that these parameters are identifiable.
@@ 3 Modeling with PDEs
where $D>0$ [L2/T] is the diffusion coefficient, and $v$ [L/T] is the advective velocity. This mathematical model is widely used in various physical, chemical and biological applications, such as the study of the dispersion of dissolved solutes (e.g. nutrients, pollutants) in porous media [Zheng2002]. We will consider the solution of (4) for the initial condition $u(x,0)=u_{b}+u_{0}$ for $|x|<h$, and $u(x,0)=u_{b}$ for $|x|>h$. We interpret $u_{b}\geq 0$ as a uniform background concentration of $u$, and we are particularly interested in the spatial spreading of $u$ that arises when an additional amount of $u$ is placed uniformly within distance $h>0$ from the origin, giving $u(x,0)=u_{b}+u_{0}$ for $|x|<h$, where $u_{0}>0$. This initial condition is plotted in Figure 3 where we have a background concentration $u_{b}=1$, and locally within the interval $-50<x<50$ we have $u(x,0)=2$. The solution of the mathematical model describes how this additional solute within $-50<x<50$ undergoes combined advection and diffusive transport as a function of position $x$ and time $t$.
@@
Given our log-likelihood function and our estimate of $\hat{\theta}$, we can work with a normalized log-likelihood function and repeat the construction of the four univariate profile likelihood functions in exactly the same way as we did for the additive Gaussian noise model. The univariate profile likelihood functions in Figure 4 indicate that the four parameters are practically identifiable. As before, we can sample the log-likelihood function to obtain $M$ parameter samples $\theta$ with $\bar{\ell}\geq\bar{\ell}^{*}$ and use these $M$ solutions of (4) to construct the 95% prediction interval given in Figure 4. For the multiplicative noise model we see that the prediction intervals have the attractive property that the upper and lower bounds are always non-negative. This is very different to working with an additive Gaussian noise model since this procedure can lead to negative prediction intervals which is not physically realistic.
@@ 5 Extensions and general remarks
In terms of making model predictions, we always used a very simple rejection sampling method to find $M$ samples of $\theta$ where the log-likelihood function is greater than some asymptotic threshold. We chose to use rejection sampling because it is both simple to implement and interpret, but other approaches are possible. For example, we could have simply used a uniform discretisation of the log-likelihood function and evaluated $\bar{\ell}$ across a uniform mesh to propagate the parameter confidence set through to examine model predictions. Both approaches carry advantages and disadvantages and give very similar results provided that $M$ is chosen to be sufficiently large when using rejection sampling, and that the uniform mesh is taken to be sufficiently dense when working with a gridded log-likelihood function.